# Advancing Prompt Learning through an External Layer

### v1

@@ 4.5. Analyzing of Multi-modal Designs
Method Base Novel HM Few-shot Task
@@ 5.1. Prompt Learning in Vision Language Models
Recently, the strong generalization capability of CLIP (Radford et al., 2021) has made it a foundation for many methods (Fang et al., 2019) in adapting pre-trained VLMs for downstream tasks (Ding et al., 2023; Cao et al., 2024b, a; Luo et al., 2023b; Sanghi et al., 2022; Li et al., 2023b, a; Pan et al., 2024; Luo et al., 2024; Di et al., 2024; Gao et al., 2023; Maaz et al., 2022; Bangalath et al., 2022; Zhang et al., 2021a; Yang et al., 2024b). Prompt learning (Liu et al., 2021; Lester et al., 2021; Li and Liang, 2021) is a widely used technique in NLP for learning downstream tasks. The use of text prompts, which are instructions provided to the language branch of a Vision-Language model (VLM), is a common practice to enhance task understanding. CoOp (Zhou et al., 2022c) and CoCoOp (Zhou et al., 2022b) fine-tune the CLIP model specifically for few-shot image recognition by optimizing a continuous set of token embeddings in the language branch. The image-conditional prompt of CoCoOp significantly contributes to enhancing generalization to unseen classes (Fang et al., 2023, 2022a, 2022b) and mitigating the risk of overfitting to the limited labeled data. Moreover, some approaches (Yao et al., 2023a; Lu et al., 2022; Zhu et al., 2023; Yao et al., 2023b) constrain the proposed learnable prompts to contain the essential general knowledge and prior distribution learning. By conditioning prompts on visual features, CoCoOp (Zhou et al., 2022b) ensures that the language model attends to relevant visual information when generating predictions. In addition to single-modal prompt tuning (Jia et al., 2022), some approaches (Khattak et al., 2023; Liu et al., 2023) introduce the multi-modal prompt-tuning designs in CLIP to effectively align V-L representations. However, their methods have not fully released the potential generalization ability of CLIP.
@@ 6. Conclusion
Prompt learning is a promising method for adapting pre-trained vision-language models. However, these methods often struggle to tackle the challenge of generalization in unseen tasks effectively. In this paper, we propose a paradigm called EnPrompt with a novel External Layer (EnLa). Specifically, we propose textual external layer and learnable visual embeddings for adapting VLMs to downstream tasks. To align the textual and visual features, we propose a novel two-pronged approach. (a) We introduce the optimal transport as the discrepancy metric to align the vision and text modalities. (b) We introduce a novel strengthening feature to enhance the interaction between these two modalities. Extensive experiments clearly demonstrated the effectiveness of our method. In the future, we will extend our effort to more challenging tasks, such as video moment retrieval (Yang et al., 2021, 2022), and investigate the cross-domain generalization ability (Yang et al., 2024a) comprehensively.

### v3

@@ 4.5. Analyzing of Multi-modal Methods
Multi-modal Methods Base Novel HM Few-shot Task
@@ 5.1. Prompt Learning in Vision Language Models
Recently, the strong generalization capability of CLIP (Radford et al., 2021) has made it a foundation for many methods (Fang et al., 2019; Cao et al., 2024b, a) in adapting pre-trained VLMs for downstream tasks (Ding et al., 2023; Luo et al., 2023b; Li et al., 2023b, a). Prompt learning (Liu et al., 2021; Lester et al., 2021; Li and Liang, 2021) is a widely used technique in NLP for learning downstream tasks. The use of text prompts, which are instructions provided to the language branch of a Vision-Language model (VLM), is a common practice to enhance task understanding (Gao et al., 2023; Zhang et al., 2021a; Yang et al., 2024b). CoOp (Zhou et al., 2022c) and CoCoOp (Zhou et al., 2022b) fine-tune the CLIP model specifically for few-shot image recognition by optimizing a continuous set of token embeddings in the language branch. The image-conditional prompt of CoCoOp significantly contributes to enhancing generalization to unseen classes (Fang et al., 2023, 2022a, 2022b) and mitigating the risk of overfitting to the limited labeled data. Moreover, some approaches (Yao et al., 2023a; Lu et al., 2022; Zhu et al., 2023; Yao et al., 2023b) constrain the proposed learnable prompts to contain the essential general knowledge and prior distribution learning. By conditioning prompts on visual features, CoCoOp (Zhou et al., 2022b) ensures that the language model attends to relevant visual information when generating predictions. In addition to single-modal prompt tuning (Jia et al., 2022), some approaches (Khattak et al., 2023; Liu et al., 2023) introduce the multi-modal prompt-tuning designs in CLIP to effectively align V-L representations (Pan et al., 2024; Luo et al., 2024; Di et al., 2024). However, their methods have not fully released the potential generalization ability of CLIP.
@@ 6. Conclusion
Prompt learning is a promising method for adapting pre-trained vision-language models. However, these methods often struggle to tackle the challenge of generalization on unseen tasks effectively. In this paper, we propose a paradigm called EnPrompt with a novel External Layer (EnLa). Specifically, we propose a textual external layer and learnable visual embeddings for adapting VLMs to downstream tasks. To align the textual and visual features, we propose a novel two-pronged approach. (a) We introduce the optimal transport as the discrepancy metric to align the vision and text modalities. (b) We introduce a novel strengthening feature to enhance the interaction between these two modalities. Extensive experiments clearly demonstrated the effectiveness of our method. In the future, we will extend our effort to more challenging tasks, such as video moment retrieval (Yang et al., 2021, 2022), and investigate the cross-domain generalization ability (Yang et al., 2024a) comprehensively.